OpenStudy (lgbasallote):

the m' for a homogeneous linear DE that has \(y = 3e^x \sin^2 x\) as particular solution is/are what? what does this question even mean?

OpenStudy (lgbasallote):

my main problem here is i dont know what m' means

OpenStudy (lgbasallote):

@UnkleRhaukus please help T_T

OpenStudy (unklerhaukus):

\[y_c=Ae^{mx}\]

OpenStudy (unklerhaukus):

but where is the DE?

OpenStudy (anonymous):

And just to clarify, m' is the first derivative of m, m'' would be the second derivative, m''' would be the third and so on...

OpenStudy (lgbasallote):

that's what i dont get... i know m is the root...so what does m' mean o.O i take the derivative of a root?

OpenStudy (unklerhaukus):

say you had a DE , like \[y''-y'-2y=0\] \[y=Ae^{mx};\qquad y'=Ame^{mx};\qquad y''=Am^2e^{mx}\] \[Am^2e^{mx}-Ame^{mx}-2Ae^{mx}=0\]\[A(m^2-m-2)e^{mx}=0\]\[m^2-m-2=0\]\[(m-2)(m+1)=0\]\[m=2,-1\] \[y=Be^{2x}+Ce^{-x}\]

OpenStudy (lgbasallote):

okay...im not getting you....

OpenStudy (unklerhaukus):

the m's are the roots of the homogenous equation

OpenStudy (lgbasallote):

i think the roots in this problem is \[m = 1 \pm i\] right?

OpenStudy (lgbasallote):

hmm how come it's sin^2 o.O

OpenStudy (unklerhaukus):

are you ment to use reduction of order because you now know a particular solution.

OpenStudy (lgbasallote):

i dont know what you're talking bout...im not good with fancy terminologies

OpenStudy (anonymous):

Do you know which satisfy the equation are called solutions ??

OpenStudy (lgbasallote):

hmm?

OpenStudy (anonymous):

Finding the particular solution mean that we are finding the solution to this DE which will satisfy this DE..

OpenStudy (lgbasallote):

perhaps a demo? im not good with words...im dyslexic with words heh

OpenStudy (anonymous):

Can I attach a file??

OpenStudy (lgbasallote):

uhmm sure?

OpenStudy (anonymous):

Firstly I should give you demo right??

OpenStudy (lgbasallote):

hmm nevermind i got unkle's demo already

OpenStudy (lgbasallote):

perhaps just a guidance how to solve this now

OpenStudy (anonymous):

This is simple one not related to m ..

OpenStudy (anonymous):

Uncle can explain you far more better than me..

OpenStudy (lgbasallote):

aww @UnkleRhaukus come back T_T

OpenStudy (unklerhaukus):

hmm

OpenStudy (wasiqss):

this is easy :D

OpenStudy (lgbasallote):

really? so how?

OpenStudy (wasiqss):

yeh man it is !

OpenStudy (lgbasallote):

so how?

OpenStudy (wasiqss):

man take 3 aside ok

OpenStudy (lgbasallote):

im not really interested whether t's hard or not..what i need to know is how

OpenStudy (lgbasallote):

it's*

OpenStudy (lgbasallote):

yeah sure...then next?

OpenStudy (wasiqss):

then open sin^2x

OpenStudy (lgbasallote):

open?

OpenStudy (wasiqss):

i mean in terms of double angle of cos

OpenStudy (wasiqss):

(1-cos2x)/2

OpenStudy (lgbasallote):

oh...still cant see it though

OpenStudy (wasiqss):

now take 1/2 as common ok?

OpenStudy (lgbasallote):

\[\frac 32 e^x(1 - \cos (2x) )?\]

OpenStudy (wasiqss):

yup

OpenStudy (lgbasallote):

so a = 1 and b = 2? \[m = 1\pm 2i?\]

OpenStudy (wasiqss):

i have a different apporach from here

OpenStudy (lgbasallote):

oh cool..i see it now..thanks

OpenStudy (wasiqss):

lgb can you do or should i help

OpenStudy (anonymous):

Go through this @lgbasallote

OpenStudy (lgbasallote):

\[\frac{3e^x}{2} - \frac{3e^x \cos 2x}{2}\] so m = 1 and m = \(1 \pm 2i\)